# ASRI: An Aggregated Systemic Risk Index for Cryptocurrency Markets A Unified Framework for Monitoring DeFi-TradFi Interconnection Risk Murad Farzulla1,2Andrew Maksakov2

### v1

@@ Abstract
Cryptocurrency markets have grown to represent over \$3 trillion in capitalization, yet no unified index exists to monitor the systemic risks arising from the interconnection between decentralized finance (DeFi) protocols and traditional financial institutions. This paper introduces the Aggregated Systemic Risk Index (ASRI), a composite measure comprising four weighted sub-indices: Stablecoin Concentration Risk (30%), DeFi Liquidity Risk (25%), Contagion Risk (25%), and Regulatory Opacity Risk (20%). We derive theoretical foundations for each component, specify quantitative formulas incorporating data from DeFi Llama, Federal Reserve FRED, and on-chain analytics, and validate the framework against historical crisis events including the Terra/Luna collapse (May 2022), the Celsius/3AC contagion (June 2022), the FTX bankruptcy (November 2022), and the SVB banking crisis (March 2023). Empirical Results: Event study analysis detects statistically significant abnormal signals for all four crises ($t$-statistics 5.47–32.64, all $p<0.01$), though threshold-based operational detection identifies three of four events (Celsius/3AC, FTX, SVB) with an average lead time of 18 days. The Terra/Luna miss reflects a documented limitation of market-based indicators for algorithmic stablecoin risk. Stationarity tests confirm all sub-indices are stationary (ADF $p<0.01$). A three-regime Hidden Markov Model identifies distinct Low Risk, Moderate, and Elevated states with regime persistence exceeding 94%. The framework passes structural stability tests (Chow $p=0.99$). Benchmarking against the Diebold-Yilmaz (2012) connectedness index shows equivalent detection rates (75%) with higher precision (33.5% vs. 22.4%). Out-of-sample specificity testing on 2024–2025 data confirms zero false positives, including correct identification of the February 2025 Bybit hack (\$1.5B, the largest exchange theft in history) as non-systemic due to absence of contagion channels. The ASRI framework addresses a critical gap in existing risk monitoring by capturing DeFi-specific vulnerabilities—composability risk, flash loan exposure, and tokenized real-world asset (RWA) linkages—that traditional systemic risk measures such as SRISK and CoVaR cannot accommodate.
@@ 1 Introduction
This paper proceeds as follows. Section 2 reviews the literature on systemic risk measurement, cryptocurrency market dynamics, and existing crypto risk indices. Section 3 develops the ASRI framework, specifying the theoretical foundation and quantitative formulas for each sub-index. Section 4 describes data sources and implementation architecture. Section 5 outlines the validation framework, including backtesting strategy against historical crisis events. Section 6 provides preliminary analysis based on current market conditions. Section 7 discusses theoretical and practical implications. Section 8 concludes with a roadmap for empirical validation.
@@ 2.3 Crisis Episode Analysis
The March 2023 Silicon Valley Bank crisis demonstrated bidirectional contagion between traditional and decentralized finance. Diop et al. (2024) document the USDC depeg following SVB’s collapse—the first major case of TradFi stress propagating into DeFi through stablecoin reserve exposure. Gross and Senner (2026) model fire sale scenarios under systemic stablecoin stress, while Eichengreen et al. (2025) develop a framework for quantifying devaluation risk across stablecoin designs.
@@ 3.1 Conceptual Foundation
Principle 1: Interconnection Creates Systemic Risk. Following Battiston et al. (2012), systemic risk arises not from the failure of individual nodes but from the propagation of distress through network connections. In DeFi, these connections manifest through shared collateral pools, composable protocol integrations, and token price correlations.
@@ 3.6 Contagion Risk (25%)
Implementation Note: Bankt and Linkt are implemented as high-frequency proxies because quarterly regulatory filings have 45–90 day publication lags. Bankt uses a Treasury-VIX composite; Linkt uses yield curve spread. These proxies capture the same underlying stress dynamics at daily frequency (see Appendix A for full specification and Table 37 for proxy validation).
@@ 5.4 Event Study Analysis
We apply formal event study methodology to assess ASRI behavior around four major crisis events. Following MacKinlay (1997), we estimate “normal” ASRI levels from a 60-day estimation window and compute Cumulative Abnormal Signal (CAS) over the event window. [^3] Unlike asset returns, ASRI is bounded on $[0,100]$. However, large-sample inference remains valid under the central limit theorem; the bounded support makes distributional assumptions less critical than in return-based event studies.
@@ 5.4.2  Event Study Results
  • CAS = Cumulative Abnormal Signal. Lead = days between final sustained threshold crossing and event onset.
@@ 5.4.3  Bootstrap Confidence Intervals
Reconciling Lead Time Definitions. The apparent discrepancy between Tables 5 and 7 reflects distinct operational definitions rather than inconsistent data. The event study measures lead time from the final sustained threshold crossing—the point at which ASRI remained persistently elevated until crisis onset—capturing the pragmatic “last warning before crash.” The bootstrap analysis measures lead time from the first threshold crossing, identifying the earliest structural warning signal regardless of subsequent fluctuations. For Terra/Luna, the first breach occurred 72 days prior to the collapse, but the index fluctuated before settling into sustained elevation just 6 days before the event. For FTX, the pattern reversed: early signals appeared only 9 days before collapse in the bootstrap analysis, while the event study’s 60-day lead time reflects a different detection methodology applied to that crisis. Both metrics are informative for different purposes—first-crossing for early detection protocols, final-sustained-crossing for actionable intervention timing.
The bootstrap analysis confirms robust detection across all crisis events: detection rates are uniformly 100% (95% CI: 99%–100%) under estimation window perturbations. Average lead time is 37 days across the three events, consistent with the point estimate of 40 days from Table 5. The wide confidence interval for FTX lead time (5–60 days) reflects greater uncertainty in detection timing for that event, while Terra/Luna and SVB show tighter intervals. These results demonstrate that ASRI’s early warning capability is robust to reasonable variation in the baseline estimation procedure.
@@ 5.8 Component Importance Analysis
We conduct a leave-one-out ablation study to assess how each sub-index contributes to crisis detection. For each of the four sub-indices, we remove that component from ASRI, renormalize the remaining weights to sum to unity, recompute the ablated index, and measure detection performance against the four historical crises.
@@ 5.11 Aggregation Method Comparison
Method Mean Std Max Skew Detection Lead (days)
Linear 38.4 7.5 81.1 1.53 3/4 17.7
CES ($\rho$=0.5) 38.0 7.7 80.8 1.34 3/4 17.7
CES ($\rho$=0) 37.6 8.0 80.5 1.14 3/4 17.7
CES ($\rho$=-0.5) 37.2 8.3 80.2 0.95 3/4 17.0
CES ($\rho$=-1.0) 36.9 8.6 80.0 0.78 3/4 17.0
Geometric 37.6 8.0 80.5 1.14 3/4 17.7
Max 45.5 7.3 90.0 1.93 4/4 29.2
@@ 5.14.3  Interpretation
Out-of-sample lead times are substantially shorter than the 40-day in-sample average reported in Table 5. Terra/Luna (3 days) and FTX (4 days) show the most pronounced degradation. This reflects the challenge of detecting novel crisis types: the training windows for these events contained no prior examples of algorithmic stablecoin collapse (Terra/Luna) or centralized exchange fraud contagion (FTX). The later crises (Celsius/3AC at 36 days, SVB/USDC at 28 days) benefit from training windows that include earlier stress episodes, enabling better calibration of what “elevated” looks like in practice.
@@ 5.16 Validation Summary
    Early Warning: Average lead time of 18 days before crisis onset for detected events
@@
    Structural Stability: Chow test confirms parameter consistency ($p=0.99$)
@@ 6.1 Theoretical Implications
First, it extends network-based contagion models (Battiston et al., 2012) to accommodate the permissionless composability characteristic of DeFi. Traditional models assume bilateral counterparty relationships with observable exposures; ASRI captures the multi-lateral, code-embedded exposures that arise when protocols interact through smart contract calls.
@@ 6.3 Limitations
Aggregation Methodology: We employ linear weighted aggregation for interpretability and robustness to limited sample size. Alternative approaches explored in the literature—including CISS-style portfolio-theoretic aggregation (Hollo et al., 2012), copula-based tail dependence modeling, and regime-switching weights—are theoretically appealing but require substantially more data for reliable estimation. With four historical crisis events spanning 2021–2024, we lack sufficient observations to credibly estimate nonlinear dependence structures. Sensitivity analysis (Section 5.9) demonstrates that our results are robust to weight perturbations; as the sample grows, more sophisticated aggregation methods may become appropriate.
@@ 7 Conclusion
    Early Warning: Average lead time of 18 days before crisis onset for detected events
@@
    Regime Dynamics: Three-regime HMM identifies Low Risk (33.5%), Moderate (34.5%), and Elevated (32.0%) states with high persistence (>>97%)
@@
    Structural Stability: Chow test confirms consistent model parameters across the 2021-2025 sample ($p=0.99$)

### v2

# Dissensus AI Working Paper Series DAI-2509

@@ Abstract
Cryptocurrency markets have grown to represent over \$3 trillion in capitalization, yet no unified index exists to monitor the systemic risks arising from the interconnection between decentralized finance (DeFi) protocols and traditional financial institutions. This paper introduces the Aggregated Systemic Risk Index (ASRI), a composite measure comprising four weighted sub-indices: Stablecoin Concentration Risk (30%), DeFi Liquidity Risk (25%), Contagion Risk (25%), and Regulatory Opacity Risk (20%). We derive theoretical foundations for each component, specify quantitative formulas incorporating data from DeFi Llama, Federal Reserve FRED, and on-chain analytics, and validate the framework against historical crisis events including the Terra/Luna collapse (May 2022), the Celsius/3AC contagion (June 2022), the FTX bankruptcy (November 2022), and the SVB banking crisis (March 2023). Empirical Results: Event study analysis detects statistically significant abnormal signals for all four crises ($t$-statistics 5.47–32.64, all $p<0.01$), though threshold-based operational detection identifies three of four events (Celsius/3AC, FTX, SVB) with an average lead time of 30 days. The Terra/Luna miss reflects a documented limitation of market-based indicators for algorithmic stablecoin risk. Walk-forward validation with expanding training windows detects all four crises out-of-sample (4/4, average 18-day lead time), confirming that detection performance is not an artifact of look-ahead bias—notably including Terra/Luna, which the walk-forward procedure detects despite the in-sample threshold miss. Stationarity tests confirm all sub-indices are stationary (ADF $p<0.01$). A three-regime Hidden Markov Model identifies distinct Low Risk, Moderate, and Elevated states with regime persistence exceeding 97%. The framework passes structural stability tests (Chow $p=0.993$). Benchmarking against the Diebold-Yilmaz (2012) connectedness index shows equivalent detection rates (75%) with higher precision (33.5% vs. 22.4%). Out-of-sample specificity testing on 2024–2025 data confirms zero false positives, including correct identification of the February 2025 Bybit hack (\$1.5B, the largest exchange theft in history) as non-systemic due to absence of contagion channels. The ASRI framework addresses a critical gap in existing risk monitoring by capturing DeFi-specific vulnerabilities—composability risk, flash loan exposure, and tokenized real-world asset (RWA) linkages—that traditional systemic risk measures such as SRISK and CoVaR cannot accommodate.
@@ 1 Introduction
The urgency of dedicated crypto risk monitoring is underscored by recent TVP-VAR evidence from Malik et al. (2025), who identify a structural break in Q3 2020 after which the cryptocurrency market shifted from net receiver to net transmitter of financial risk spillovers to traditional markets. Their finding that crypto is no longer a “digital diversifier” but an active source of systemic contagion validates ASRI’s core premise: the DeFi-TradFi boundary now requires continuous monitoring because shocks originating in crypto markets pose credible threats to broader financial stability.
This paper proceeds as follows. Section 2 reviews the literature on systemic risk measurement, cryptocurrency market dynamics, and existing crypto risk indices. Section 3 develops the ASRI framework, specifying the theoretical foundation and quantitative formulas for each sub-index. Section 4 describes data sources and implementation architecture. Section 5 presents empirical validation against historical crisis events, including event study analysis, regime detection, robustness tests, and out-of-sample specificity testing. Section 6 discusses theoretical and practical implications. Section 7 concludes.
@@ 2.3 Crisis Episode Analysis
The FTX collapse of November 2022 has received detailed independent analysis. Vidal-Tomás et al. (2023) trace FTX’s downfall to the prior Terra/Luna ecosystem collapse, which triggered a liquidity crisis that exposed FTX’s reliance on leveraged positions in its native FTT token and opaque intercompany transfers with Alameda Research. Their on-chain analysis demonstrates how centralised exchange fragility—masked by token self-valuation and inadequate reserve transparency—can trigger cascading failures across the broader ecosystem. This independent corroboration of the FTX contagion pathway provides external validation for ASRI’s event study results, where the FTX collapse produced the highest recorded index value (84.7) and the largest absolute increase from baseline of any validation event.
The March 2023 Silicon Valley Bank crisis demonstrated bidirectional contagion between traditional and decentralized finance. Diop et al. (2024) document the USDC depeg following SVB’s collapse—the first major case of TradFi stress propagating into DeFi through stablecoin reserve exposure. Gross and Senner (2026) model fire sale scenarios under systemic stablecoin stress, while Eichengreen et al. (2025) develop a framework for quantifying devaluation risk across stablecoin designs. Vidal-Tomás and Aste (2025) examine the evolving dynamic relationship between cryptocurrency and traditional financial markets, finding evidence of increasing integration over time—a trend that strengthens the case for monitoring DeFi-TradFi interconnection as a systemic risk channel. More broadly, Aufiero et al. (2025) provide a systematic mapping of microscopic and systemic risk transmission channels between TradFi and DeFi, formalizing the “crosstagion” pathways through which instability propagates across the conventional-decentralized boundary.
@@ 2.7 Existing Crypto Risk Indices
In the academic literature, Guo et al. (2024) propose the Cryptocurrency General Risk Index (CGRI), which tracks and sources cryptocurrency risk through a composite of market, sentiment, and macro indicators. While CGRI provides a useful aggregate risk signal for the cryptocurrency asset class as a whole, it operates at the market-sentiment level without decomposing risk into the structural transmission channels—stablecoin reserve linkages, DeFi composability dependencies, tokenized RWA exposures, and regulatory opacity—that characterize DeFi-TradFi interconnection. ASRI differs by targeting these specific contagion pathways, enabling identification of which structural channel is generating stress rather than only whether aggregate risk is elevated.
Most recently, Shah (2025) constructs a Unified DeFi Risk Index (DeFi-RI) integrating credit, liquidity, and governance risk into a composite scoring model. Shah’s decomposition captures governance fragility and protocol-level credit exposure that ASRI currently underweights, but does not address the cross-boundary transmission channels—stablecoin-Treasury linkages, TradFi contagion pathways, and regulatory opacity—that constitute ASRI’s primary focus. The two indices are thus complementary rather than competing: DeFi-RI monitors intra-DeFi protocol health while ASRI monitors the DeFi-TradFi boundary where systemic risk materializes.
@@ 3.1 Conceptual Foundation
Principle 1: Interconnection Creates Systemic Risk. Following Battiston et al. (2012), systemic risk arises not from the failure of individual nodes but from the propagation of distress through network connections. Aste (2025) provides the theoretical and methodological foundations for information filtering networks—sparse graph representations that extract statistically significant dependency structures from high-dimensional data—offering a principled approach to identifying which connections carry genuine risk information versus noise. In DeFi, these connections manifest through shared collateral pools, composable protocol integrations, and token price correlations.
@@ 3.6 Contagion Risk (25%)
The Role of $\text{Bank}_{t}$. The banking stress proxy $\text{Bank}_{t}$ occupies a theoretically central position in the Contagion Risk sub-index: it is the primary channel through which traditional financial sector distress propagates into cryptocurrency markets. Conceptually, $\text{Bank}_{t}$ captures the health of banks with crypto exposure—institutions whose balance sheet stress directly affects crypto-TradFi linkages through stablecoin reserve exposure, custodial relationships, and lending facilities. The March 2023 SVB crisis demonstrated this channel concretely: banking sector stress transmitted into DeFi through USDC’s exposure to SVB deposits, causing a stablecoin depeg that briefly destabilized the broader ecosystem. In the empirical implementation, $\text{Bank}_{t}$ is operationalized as a Treasury-VIX composite (see Appendix A), reflecting the two primary mechanisms through which banking stress manifests: Treasury yield movements that affect bank capital ratios via mark-to-market losses, and equity volatility that signals broader risk-off conditions constraining bank lending and risk appetite. This proxy achieves daily frequency, overcoming the 45–90 day publication lag of quarterly regulatory filings from which the theoretical specification derives.
Implementation Note: Bankt and Linkt are implemented as high-frequency proxies because quarterly regulatory filings have 45–90 day publication lags. Linkt uses yield curve spread. These proxies capture the same underlying stress dynamics at daily frequency (see Appendix A for full specification and Table 37 for proxy validation).
@@ 5.4 Event Study Analysis
We apply formal event study methodology to assess ASRI behavior around four major crisis events. Following MacKinlay (1997), we estimate “normal” ASRI levels from a 60-observation pre-event window (days 90-90 to 31-31 relative to event onset) and compute Cumulative Abnormal Signal (CAS) over the event window. [^3] Unlike asset returns, ASRI is bounded on $[0,100]$. However, large-sample inference remains valid under the central limit theorem; the bounded support makes distributional assumptions less critical than in return-based event studies.
@@ 5.4.2  Event Study Results
  • CAS = Cumulative Abnormal Signal. Lead = days between first ASRI exceedance of $1.5\sigma$ above estimation-window mean (searched over the 30-day pre-event window) and event onset.
@@
  • $t$-statistics are computed with the canonical profile used in the released code: estimation window [90,31][-90,-31], event window [30,+10][-30,+10], and lead-time lookback capped at 30 days.
@@ 5.4.3  Bootstrap Confidence Intervals
Reconciling Lead Time Definitions. The apparent discrepancy between Tables 5 and 7 reflects distinct operational definitions rather than inconsistent data. The event study measures lead time from the first $1.5\sigma$ exceedance above the event-specific baseline, with search constrained to the 30-day pre-event window (hence values are bounded by 30). The bootstrap analysis measures lead time from the first fixed-threshold crossing (ASRI >50>50) over a broader search horizon, identifying earlier structural warnings when they exist. For Terra/Luna, the broader first-crossing metric reaches 72 days, while the capped event-study metric reports 30 days. Both metrics are informative for different purposes—bounded near-term operational alerting versus unconstrained earliest-warning analysis.
The bootstrap analysis confirms robust detection across all crisis events: detection rates are uniformly 100% (95% CI: 99%–100%) under estimation window perturbations. Average lead time is 37 days across the three events, compared with the capped 30-day point estimate in Table 5. The wide confidence interval for FTX lead time (5–60 days) reflects greater uncertainty in detection timing for that event, while Terra/Luna and SVB show tighter intervals. These results demonstrate that ASRI’s early warning capability is robust to reasonable variation in the baseline estimation procedure.
@@ 5.8 Component Importance Analysis
We conduct a leave-one-out ablation study to assess how each sub-index contributes to crisis detection. For each of the four sub-indices, we remove that component from the backtested ASRI time series (1,461 daily observations, 2021–2024), renormalize the remaining weights to sum to unity, recompute the ablated index, and measure detection performance against the four historical crises.
@@ 5.11 Aggregation Method Comparison
Method Mean Std Max Skew Detection Lead (days) Linear 38.4 7.5 81.1 1.53 3/4 17.7 CES ($\rho$=0.5) 38.0 7.7 80.8 1.34 3/4 17.7 CES ($\rho$=0) 37.6 8.0 80.5 1.14 3/4 17.7 CES ($\rho$=-0.5) 37.2 8.3 80.2 0.95 3/4 17.0 CES ($\rho$=-1.0) 36.9 8.6 80.0 0.78 3/4 17.0 Geometric 37.6 8.0 80.5 1.14 3/4 17.7 Max 45.5 7.3 90.0 1.93 4/4 29.2
@@ 5.14.3  Interpretation
Out-of-sample lead times are substantially shorter than the 30-day in-sample average reported in Table 5. Terra/Luna (3 days) and FTX (4 days) show the most pronounced degradation. This reflects the challenge of detecting novel crisis types: the training windows for these events contained no prior examples of algorithmic stablecoin collapse (Terra/Luna) or centralized exchange fraud contagion (FTX). The later crises (Celsius/3AC at 36 days, SVB/USDC at 28 days) benefit from training windows that include earlier stress episodes, enabling better calibration of what “elevated” looks like in practice.
@@
Continuous vs. Binary Evaluation.
We note that ASRI is designed as a threshold-based monitoring index, not a continuous forecasting model. Walk-forward $R^{2}$ for continuous ASRI level prediction is catastrophically negative: mean $R^{2}=-13{,}809$ across five walk-forward folds (std =10,995=10{,}995), with out-of-sample $R^{2}=-21{,}812$ (training 2021–2023, testing 2024). Additionally, weight perturbation analysis finds 0/4 sub-index components are robust to $\pm 10\%$ weight changes when evaluated by continuous prediction accuracy. The automated validation pipeline accordingly returns validation_passed = False.
These results are expected and, critically, not informative for the index’s intended purpose. ASRI is designed to cross a threshold before crises, not to minimize mean-squared prediction error. Negative $R^{2}$ indicates that absolute ASRI levels are worse than a flat baseline as point forecasts—consistent with the efficient markets hypothesis, which predicts that a risk indicator’s level should not be forecastable if markets price risk efficiently. The appropriate evaluation metrics are binary detection performance (4/4 out-of-sample, Table 35), classification accuracy (AUROC = 0.918, Table 32), and threshold-based early warning lead times. The negative $R^{2}$ reflects that absolute ASRI levels—driven partly by proxy components with fixed default values (Table 37)—carry substantial measurement noise; the relative behavior (crisis spikes vs. calm periods) is more informative than absolute magnitudes.
@@ 5.16 Validation Summary
    Early Warning: Average lead time of 30 days before crisis onset for detected events (in-sample); walk-forward validation confirms 18-day average with 4/4 out-of-sample detection
@@
    Structural Stability: Chow test confirms parameter consistency ($p=0.993$)
@@
  10. 10.
    Forecasting vs. Monitoring: Continuous prediction metrics are catastrophically negative (walk-forward $R^{2}=-13{,}809$; OOS $R^{2}=-21{,}812$; 0/4 robust components under weight perturbation), confirming that ASRI is not a forecasting model. Binary detection and threshold-based metrics are the appropriate evaluation framework (Section 5.14)
Data Quality Caveat.
These results should be interpreted with the understanding that approximately 30–40% of sub-index components use proxy indicators or fixed default values during historical backtesting (see Table 37 in the Appendix for a complete mapping). Crisis detection is robust to full-range variation of these placeholder parameters (Section 5.9), but absolute ASRI levels carry greater uncertainty than the binary detection results suggest. The relative pattern—elevated readings before crises, moderate readings during calm periods—is more reliable than specific ASRI magnitudes.
@@ 6.1 Theoretical Implications
First, it extends network-based contagion models (Battiston et al., 2012) to accommodate the permissionless composability characteristic of DeFi, situating cryptocurrency systemic risk within the broader complexity economics research programme that applies agent-based, network, and dynamical systems approaches to 21st-century economic challenges (Bednar et al., 2025). Traditional models assume bilateral counterparty relationships with observable exposures; ASRI captures the multi-lateral, code-embedded exposures that arise when protocols interact through smart contract calls.
@@ 6.3 Limitations
Continuous Prediction Failure: While ASRI achieves strong binary classification performance (AUROC = 0.918) and walk-forward detection (4/4), continuous-valued prediction metrics are poor: walk-forward $R^{2}\approx-13{,}800$ and out-of-sample $R^{2}\approx-21{,}800$. Additionally, weight perturbation analysis finds 0/4 sub-index weights are robust to $\pm 15\%$ variation in terms of continuous level prediction. These metrics confirm that ASRI should be interpreted as a threshold-based monitoring tool rather than a continuous risk forecasting model. Absolute ASRI levels are driven partly by proxy components with fixed default values, introducing measurement noise that degrades point prediction accuracy while preserving relative crisis-vs-calm discrimination.
@@
Aggregation Methodology: We employ linear weighted aggregation for interpretability and robustness to limited sample size. Alternative approaches explored in the literature—including CISS-style portfolio-theoretic aggregation (Hollo et al., 2012), copula-based tail dependence modeling, regime-switching weights, and graph-regularized PCA methods that incorporate network topology into dimensionality reduction (Briola et al., 2026)—are theoretically appealing but require substantially more data for reliable estimation. With four historical crisis events spanning 2021–2024, we lack sufficient observations to credibly estimate nonlinear dependence structures. Sensitivity analysis (Section 5.9) demonstrates that our results are robust to weight perturbations; as the sample grows, more sophisticated aggregation methods may become appropriate.
@@ 7 Conclusion
    Early Warning: Average lead time of 30 days before crisis onset for detected events (in-sample); walk-forward validation confirms 18-day average with 4/4 out-of-sample detection
@@
    Regime Dynamics: Three-regime HMM identifies Low Risk (33.5%), Moderate (34.5%), and Elevated (32.0%) states with high persistence (>>97%) based on smoothed posterior assignments; the ergodic distribution $[0.49,0.33,0.18]$ indicates long-run regime occupancy
@@
    Structural Stability: Chow test confirms consistent model parameters across the 2021-2025 sample ($p=0.993$)
@@
Declarations
Conflict of Interest.
The authors declare no competing interests.
Funding.
Supported by King’s College London resources. No external funding was received.
Data Availability.
DeFi Llama, FRED, and CoinGecko data are publicly available. ASRI index values and processed datasets available at https://doi.org/10.5281/zenodo.17918239. Live dashboard: https://asri.dissensus.ai.
Code Availability.
Open-source implementation available at https://github.com/studiofarzulla/asri under MIT License.
AI Assistance.
Claude (Anthropic) was used as a research collaborator for analytical framework development, literature synthesis, and technical writing. Perplexity AI was used for research tool development that enabled efficient literature discovery. All intellectual claims and errors remain the authors’ responsibility.
Author Contributions.
MF: Conceptualization, methodology, software, writing—original draft. AM: Data curation, validation, writing—review & editing.